OpenStudy (king):

Help............in the decimal number 33xy2y ,x and y represent digits. If the number is divisible by 275 find x. Help............in the decimal number 33xy2y ,x and y represent digits. If the number is divisible by 275 find x. @Mathematics

OpenStudy (anonymous):

338525

OpenStudy (king):

how?

OpenStudy (anonymous):

x = 8 y = 5

OpenStudy (anonymous):

see as it is divisible by 275 means y will be either 5 or, 0

OpenStudy (king):

yeah kk

OpenStudy (anonymous):

and 275 = 5*5*11

OpenStudy (king):

yes

OpenStudy (anonymous):

that is the resultant should be divisible by 11

OpenStudy (king):

yeah

OpenStudy (anonymous):

thus the 33xy2y should be divisible by 11. That is the difference of sum of digits at odd place with the sum of digits at even places should be either 0 or, 11, or, multiple of 11

OpenStudy (anonymous):

sum of odd place digits is 5+x and sum of even place digit is 3 +2y

OpenStudy (anonymous):

Difference is (5+x) - (3+2y) it should be either 0 / 11/ multiple of 11

OpenStudy (anonymous):

difference is 2 + x - 2y

OpenStudy (king):

yes

OpenStudy (anonymous):

y can be either 0 or, 5

OpenStudy (anonymous):

when y = 0 therefore we have 2 + x (it should be 0 / 11 / multiple of 11) but in no case it is possible so Y cannot be 0

OpenStudy (king):

possible 339020

OpenStudy (king):

x=9 y=0

OpenStudy (anonymous):

so when y = 5 then we have 2 + x - 2*5 = 2 + x -10 = -8 + x

OpenStudy (anonymous):

yes but then it is not divisible by 275

OpenStudy (anonymous):

if x = 9 in that case it is not divisible by 275

OpenStudy (king):

oh

OpenStudy (anonymous):

so it cannot be 9

OpenStudy (anonymous):

now we got -8 + x for what value of x u will have either 0 / 11 / multiple of 11

OpenStudy (king):

ok thnx u always make me understand thnx a lot...............!!!!!!!!

OpenStudy (anonymous):

so got the point how many rules u have to apply.........

OpenStudy (king):

yeah

OpenStudy (anonymous):

another thing that u can do is use short cut that is calculator......hahaha satisfied with the answer

OpenStudy (king):

yup 100% satisfied

OpenStudy (anonymous):

hey actually u tried to understand the thing other pplare just interested in answers not in the procedure so i gave u a medal